Bluto is trying to run away with Olive Oyl while Popeye is trying to hold onto her. Bluto pulls towards the west with 300 N of force. Popeye pulls eastward with a strength of 275 N. Olive Oyl helps Popeye with 50 N of force.
A) What is the net force acting on Olive Oyl?
B) What is Olive Oyl's acceleration if her mass is 30 kg (~66 lbs)?
C) If nobody lets go or gets tired, how long will it take to drag Olive Oyl onto Popeye's boat (5m away) where they can sail away from Bluto?
Answer #1

here,

the force exerted by Bluto , F1 = - 300 N

force exerted by Popeye , F2 = 275 N

force exerted by Olive Oyl , F3 = 50 N

a)

the net force on Olive Oyl ,Fn = F1 + F2 + F3

Fn = - 300 + 275 + 50 N = 25 N

b)

mass of Olive Oyl , m = 30 kg

the acceleration of Olive Oyl , a = Fn /m

a = 25 /30 m/s^2 = 0.83 m/s^2

c)

let the time taken be t

for s = 5m

s = u * t + 0.5 * a * t^2

5 = 0 + 0.5 * 0.83 * t^2

solving for t

t = 3.46 s

the time taken is 3.46 s